# LED: LLM Enhanced Open-Vocabulary Object Detection without Human Curated Data Generation

### v1

@@ 4.2.2 Adapter Architecture Analysis
In MLLMs, the alignment between the vision encoder and LLMs directly determines the quality of the LLMs’ understanding of image tokens. To maximize the quality of the adaptation prompts provided by LLMs, we employ InternVL ViT and its pre-trained weights as the vision encoder for models such as InternVL2-1B, 2B, and 8B. Although the hidden states from the LLM decoder contain rich semantic understanding of the images as endowed by the LLMs, they also include a significant amount of redundant information that has already been captured by the vision encoder of the detector. In this section, we discuss various design approaches to extract adaptation prompts from the hidden states, as well as how to handle the remaining text embeddings resulting from the truncation of the hidden states.
We analyze and evaluate an optimal feature fusion strategy to design the adapter, we discussed various possible adapter design schemes:
Architecture $\mathrm{I}$ (Double Cross-Modal Fusion): As show in the Figure 6 (in the supplement), owing to the robust cross-modal feature fusion capabilities of MLLMs, the text embeddings $\mathbf{E_{T}}$ extracted from the hidden state, encompass substantial information relevant to image understanding. Furthermore, these embeddings can guide the Detector by leveraging user-provided instructions. The LLM input is formulated as a structured prompt that integrates task instructions and object captions, adhering to GroundingDINO’s annotation protocol:
<<<user>>>: [Instruction: Detect objects in this image. If present, locate them with bounding boxes.] + [Caption: {cat, dog, car,…}]‘
We extract image/text embeddings $\mathbf{E_{T}}/\mathbf{E}_{V^{D}}$ from the decoder’s hidden states using positional segmentation. The cross-attention mechanism operates through two stages:
  1. 1.
    Text-guided image fusion: Image embeddings as queries attend to text embeddings (keys/values)
  2. 2.
    Prompt-enhanced detection: Detector queries attend to fused embeddings (keys/values)
@@
$\mathbf{A_{P}}=Gatng(CrossAtt(CrossAtt(\mathbf{E_{V}}^{L},\mathbf{E_{T}}),\mathbf{E}_{V^{D}}))$ (7)
Architecture $\mathrm{II}$ (Late Prompt Projection):We removing the cross-modal attention layers to Image embedding of the Detector. As show in the Figure 7 (See supplement), A $3\!\times\!3$ convolution (stride=2) directly process the MLLM’s fusion adaptation prompts $\mathbf{A_{P}}$to match the detector cross-modal decoder’s input space, implementing:
$\mathbf{A_{P}}=\mathrm{Conv2D}(CrossAtt(\mathbf{E_{V}}^{L},\mathbf{E_{T}}))\in\mathbb{R}^{B\times L\times d}$ (8)
Then the adaptation prompts provide to the last cross-modal decoder layer $\ell_{D}=6$, as the describle in subsection 3.4.
Architecture $\mathrm{III}$ (Early Decoder Injection): Instead of feeding prompts to the detector’s final decoder layer (Arch. II), we inject them at the first decoder layer by choosing $\ell_{D}=1$ instead of $\ell_{D}=6$ as show in the Figure 7 (See supplement). This allows gradual refinement through 6 decoder blocks.
@@
Architecture $\mathrm{IV}$ (Text-Free Adaptation): As shown in the Figure 4, Eliminating text embeddings $\mathbf{E_{T}}$, we directly project image embeddings $\mathbf{E}_{v}$ using the same convolutional layer as Arch. II. This tests the hypothesis that visual embeddings alone carry sufficient multimodal signals.
Comparing Arch. I and II, adaptation prompts injection into the cross-modal decoder (II) significantly improves descr-AP by 4.17% compared to I’s 0.86%, but reduces categ-AP by 2.20%. This suggests that input fusion in the cross-modal decoder preserves categorical priors, while internal fusion enhances open-vocabulary adaptation.
Arch. I and II show similar performance on the overall task, but Arch. III’s first-layer injection achieves higher categ-AP (32.07% vs. II’s 31.05%) but lower descr-AP (19.99% vs. 20.26%). We attribute this to early fusion, which enables gradual feature refinement through decoder layers, better preservation of categorical description features, and reduced overfitting to textual descriptions. Despite the rich textual understanding of MLLM (Arch. I), Arch. IV achieves superior performance (27.87% vs. 22.57% AP). This reveals two phenomena: (1) visual embeddings $\mathbf{E}_{v}$ already encode cross-modal alignment through MLLM pretraining, and (2) naive text-visual fusion may introduce semantic noise without explicit contrastive training for multimodal alignment.
Method AP AP AP
descr,categ categ descr
@@
I 22.57% $\uparrow$0.8833.80% $\uparrow$0.5516.95% $\uparrow$0.86%
II 24.52% $\uparrow$2.8431.05% $\downarrow$2.2020.26% $\uparrow$4.17%
III 24.63% $\uparrow$2.9532.07% $\downarrow$1.1919.99% $\uparrow$3.90%
IV 25.98% $\uparrow$4.2933.89% $\uparrow$0.6321.06% $\uparrow$4.97%
@@ 4.2.3 Different LLMs
OmniLabel AP AP AP AP
descr,categ categ descr descr-pos
@@
InternVL2-1B 27.90% $\uparrow$6.2233.93% $\uparrow$0.6723.70% $\uparrow$7.6136.24% $\uparrow$11.63%
InternVL2-2B 27.87% $\uparrow$6.1933.44% $\uparrow$0.1923.90% $\uparrow$7.8136.13% $\uparrow$11.53%
InternVL2-8B 26.33% $\uparrow$4.6532.03% $\downarrow$1.2222.36% $\uparrow$6.2734.34% $\uparrow$9.73%
Table 4: OmniLabel evaluation on GroundingDINO vs. got adaptation promps InternVL-1B, InternVL-2B and InternVL-8B.
@@
L
ED
achieves significant improvements in open-set object detection, with 6.22% and 6.19% AP gains for InternVL-1B and InternVL-2B, respectively. The attention mechanisms in InternVL’s decoder layers effectively map visual features to semantically rich textual embeddings, enabling better grounding of complex object descriptions. This addresses the limitations of current mainstream detectors in aligning textual descriptions with images.
In particular, the description and positioning joint task shows remarkable 11.63% and 11.53% AP improvements, demonstrating MLLMs’ capacity for spatial relationship understanding. We hypothesize that this stems from the model’s strong logical reasoning capabilities, which enhance the understanding of spatial relationships in textual descriptions, such as ”a person holding a phone.” The visual-textual co-attention mechanisms likely help disambiguate spatial configurations that challenge pure visual encoders. This phenomenon can be clearly observed in the analysis of subsection 6.7 (See supplement).
@@
It is worth noting that category detection shows minimal improvements (0.67% AP and 0.19% gain) across experiments because closed-set category recognition relies primarily on visual prototype matching rather than semantic understanding. Thus, the Swin-Tiny encoder’s features may already saturate performance on this conventional detection task, leaving limited room for text-guided enhancement.
@@
We evaluated our model on RefCOCO/+/g [39], and the results are shown in Table 5. Combining the evaluation results from OmniLabel, we found that: InternVL-1B shows comparable performance to InternVL-8B despite 14×14$\times$14 × more parameters. This suggests a pattern of diminishing returns in scaling LLMs for visual grounding tasks, supporting the hypothesis from [7] that small model can suffice for certain vision-language mappings.
4.3 Ablations
To quantify the different-level features contained in the hidden states obtained from various decoder layers of the MLLM, and use only the vision transformer(ViT) embeddings obtained from the MLLM’s vision encoder without the LLM ($-1_{th}$ decoder layers) to provide adaptation prompts for the detector decoder, we conducted the following ablation experiments.
Figure 5: The OmniLabel evaluation assesses adaptation prompts derived from different decoder layers in the LLM, spanning layers $-1$ to $24$. The $-1$ layer indicates that adaptation prompts are generated solely using the vision encoder in the MLLM, rather than being truncated from the hidden state.
The OmniLabel evaluation for adaptation prompts from Different Decoder Layer $\ell_{LM}$ as show in the Figure 5. From the experimental results, it is evident that the additional information covered in the adaptation prompts provided by the ViT embedding of the MLLM ($-1_{th}$ layer) alone is extremely limited and even detrimental to the performance of the categories task. However, the semantic information contained in the lower-level LLM decoder remains highly rich. In the case of InternVL2-1B, which uses a 24-layer LLM Qwen2-0.5B, utilizing just 2 layers of the LLM significantly improves the detector’s performance, peaking at the fourth layer. Notably, the adaptation prompts extracted from deeper layers of the detector do not necessarily imply higher-quality feature information. This inference fully aligns with the conclusions we derived from the attention scores in the Decoder’s hidden layers in section 3 and Figure 9 (See supplement), namely that image processing is completed in the early layers, and most of the image information has already been ”condensed” into the representations.
@@ 5 Conclusions
Our work presents a systematic framework leveraging shallow MLLM decoder layers to enhance visual grounding efficiency. Key contributions include: (1) Demonstrating that early Transformer layer hidden states retain sufficient spatial-semantic correlations for grounding; (2) Proposing a zero-initialized cross-attention adapter for effective MLLM-to-detector knowledge transfer; (3) Introducing two adaptation prompt schemes that boost descriptive target precision without degrading category accuracy. The Qwen2-0.5B with Swin-Tiny as a vision encoder scheme achieves a 2.33% OmniLabel recognition improvement (vs. GroundingDINO’s 24.01%) with only 8.7% added GFLOPs, while Qwen2-0.5B with InternVL ViT yields a 6.22% gain, still improve significant on RefCOCO/+/g. Evaluations of adapter architectures, LLM scales, and detector layers validate robustness, especially in terms of describing targets and zero-shot tasks, the improvement is significant. Our generalizable adapter design enables real-time vision-language applications while mitigating overfitting risks from manual prompt bias. This advances practical MLLM deployment in downstream tasks.
Supplementary Material
6 More Implementation Details
6.1 Adapter Architecture Detail
@@
6.2 Expertiment Configuration
In subsubsection 4.2.1, the experimental parameters used in stages 1, 2, and 3 for deploying our method are shown in the Table 6. In subsubsection 4.2.3, we explore how varying the parameter sizes of LLMs affect the quality of the adaptation prompt, following the experimental setup described in Table 7. In subsubsection 4.2.2, We used the Expertiment Configuration in Table 8 to test how LLM influence the quality of adaptation prompt. OmniLabel evaluation for different adapter architectures shows on Table 3. Additionally, in subsection 4.3, we evaluate the effectiveness of adaptation prompts generated by different layers of the LLM decoder, using the configuration specified in Table 9.
Expertiment config for sharing vision decoder
@@ 6.2 Expertiment Configuration
$\mathbf{HD}$ from $\ell_{LM}$ 2
$\mathbf{AP}$ to $\ell_{D}$ 1
Training Dataset Object365, COCO, Flickr30k
@@
adapter architecture IV
Table 6: Expertiment configuration details for vision decoder share from detector.
Expertiment config for different LLMs
@@
Training Dataset Object365, COCO, Flickr30k
adapter architecture IV
Table 7: Expertiment configuration details for different LLMs.
Expertiment config for adapter architecture
@@
Training Dataset COCO, Flickr30k
Table 8: Expertiment configuration details for different adapter architecture.
Expertiment config for Ablations
@@
Training Dataset Object365, COCO, Flickr30k
adapter architecture IV
Table 9: Ablations expertiment configuration details for different decoder layer.
6.3 Dynamic Image Processing
@@
Figure 8: Dynamic Image Processing
6.4 MLLM Alignment (Swin-T & Qwen2-0.5B)
@@
6.5 Attention Maps in Qwen2-0.5B Decoding
Figure 9: Attention maps during the decoding process of Qwen2-0.5B.
Attention $\text{Attention Score}(Q,K)=\frac{QK^{T}}{\sqrt{d_{k}}}$ (10)
Where, $Q,K$ are the Query and Key matrix in the transformer Decoder layer, $\text{softmax}(\cdot)$as show in Figure 9, we have reached an approximate conclusion [3]. We can observe that in the first layer, attention is distributed relatively smoothly across different types of tokens. In the deeper layers, starting from local attention, attention scores are aggregated onto system prompts, instructions, and output tokens, while attention to image tokens becomes quite sparse. In the deeper layers, there are strong vertical lines (in the system prompts) that dominate most of the attention scores. The presence of these strong vertical lines indicates that certain input tokens remain highly attended to throughout the decoding process. If large-scale, high-intensity attention to image tokens were also maintained in the deeper layers, it would suggest that a significant amount of visual information is still needed for inference at those stages. However, as seen in the visualization, the deeper layers primarily focus on some key text tokens, which precisely indicates that the processing of images has already been completed in the earlier layers, and most of the image information has been ”condensed” into the representations.
6.6 Dataset Introduction
Table 11 provides the attributes, sizes, and functional descriptions of all datasets used in our stages 1, 2, 3 and evaluation, to illustrate the support they offer during the process.

### v2

@@ 4.2.2 Adapter Architecture Analysis
In MLLMs, how the vision encoder aligns with the LLM decisively affects image-token understanding. We adopt the InternVL ViT backbone (pre-trained) for InternVL2-1B/2B/8B and compare four adapter designs that extract adaptation prompts from decoder hidden states while coping with truncated text embeddings.
Arch. I — Double Cross-Modal Fusion. As shown in the Figure 7 (in the supplement), owing to the robust cross-modal feature fusion capabilities of MLLMs, the text embeddings $\mathbf{E_{T}}$ extracted from the hidden state, encompass substantial information relevant to image understanding. Furthermore, these embeddings can guide the Detector by leveraging user-provided instructions. The LLM input is formulated as a structured prompt that integrates task instructions and object captions, adhering to GroundingDINO’s annotation protocol:
$<$user$>$: [Instruction: ’Detect objects in this image. If present, locate them with bounding boxes.’] + [Caption: ’The cat on the right’]
We extract image/text embeddings $\mathbf{E_{T}}/\mathbf{E}_{V^{D}}$ from the decoder’s hidden states using positional segmentation. The cross-attention mechanism operates through two stages: (1) Text-guided image fusion: Image embeddings as queries attend to text embeddings (keys/values); (2) Prompt-enhanced detection: Detector queries attend to fused embeddings (keys/values)
@@
$\mathbf{A_{P}}=Gating(CrossAtt(CrossAtt(\mathbf{E_{V}^{L}},\mathbf{E_{T}}),\mathbf{E}_{V^{D}}))$ (7)
Arch. II — Late Prompt Projection. We drop the second cross-modal block. As shown in the Figure 8 (see Supplement), a $3{\times}3$ /2 convolution maps the fused prompt to the detector space and injects it into the last decoder layer ($\ell_{D}{=}6$):
$\mathbf{A_{P}}=\mathrm{Conv2D}\bigl{(}\mathrm{CrossAtt}(\mathbf{E_{V}}^{L},\mathbf{E_{T}})\bigr{)}\in\mathbb{R}^{B\times L\times d}$ (8)
Then the adaptation prompts provide to the last cross-modal decoder layer $\ell_{D}=6$, as the describle in Section 3.4.
Arch. III — Early Decoder Injection. Same as Arch. II but insert at the first decoder layer ($\ell_{D}{=}1$) for progressive refinement:
@@
Arch. IV — Text-Free Adaptation. As shown in the Figure 4, Eliminating text embeddings $\mathbf{E_{T}}$, we directly project image embeddings $\mathbf{E}_{v}$ using the same convolutional layer as Arch. II. This tests the hypothesis that visual embeddings alone carry sufficient multimodal signals.
As illustrated in the Table 3, Arch. II boosts descr-AP by +4.17% over Arch. I but sacrifices 2.20% categ-AP, implying that late fusion favours free-form phrases at the cost of categorical priors. Early injection (Arch. III) recovers part of the categ-AP while retaining most of the descr-AP. Surprisingly, text-free Arch. IV outperforms all others, suggesting (i) vision embeddings already encode rich cross-modal cues from pretraining, and (ii) naive text–vision fusion may introduce semantic noise without explicit contrastive alignment.
Method AP descr,categ AP categ AP descr
@@
Arch. I 22.57% $\uparrow$0.8833.80% $\uparrow$0.5516.95% $\uparrow$0.86%
Arch. II 24.52% $\uparrow$2.8431.05% $\downarrow$2.2020.26% $\uparrow$4.17%
Arch. III 24.63% $\uparrow$2.9532.07% $\downarrow$1.1919.99% $\uparrow$3.90%
Arch. IV 25.98% $\uparrow$4.2933.89% $\uparrow$0.6321.06% $\uparrow$4.97%
@@ 4.2.3 Different LLMs
L
ED
achieves significant improvements in open-set object detection, with 6.22% and 6.19% AP gains for L ED  Qwen2-0.5B and L ED  +InternLM2-1.8B, respectively. The attention mechanisms in LLM’s decoder layers effectively fusing semantically knowledge to visual features, enabling better grounding of complex object descriptions.
In particular, the description and positioning joint task shows remarkable 11.63% and 11.53% AP improvements, demonstrating MLLMs’ capacity for spatial relationship understanding. We hypothesize that this stems from the model’s strong logical reasoning capabilities, which enhance the understanding of spatial relationships in textual descriptions, such as "a person holding a phone." The visual-textual co-attention mechanisms likely help disambiguate spatial configurations that challenge pure visual encoders. This phenomenon is clearly observed in the analysis of Appendix C (see Supplement).
OmniLabel AP descr,categ AP categ AP descr AP descr-pos
@@
Qwen2-0.5B 27.90% $\uparrow$6.2233.93% $\uparrow$0.6723.70% $\uparrow$7.6136.24% $\uparrow$11.63%
InternLM2-1.8B 27.87% $\uparrow$6.1933.44% $\uparrow$0.1923.90% $\uparrow$7.8136.13% $\uparrow$11.53%
InternLM2.5-7B 26.33% $\uparrow$4.6532.03% $\downarrow$1.2222.36% $\uparrow$6.2734.34% $\uparrow$9.73%
Table 4: OmniLabel evaluation on GroundingDINO vs. adaptation prompts L ED  + Qwen2-0.5B, L ED  + InternLM2-1.8B and L ED  + InternLM2.5-7B.
@@
Notably, category detection shows minimal improvements (0.67% AP and 0.19% gain) across experiments because closed-set category recognition relies primarily on visual prototype matching rather than semantic understanding. Thus, the Swin-Tiny encoder’s features may already saturate performance on this conventional detection task, leaving limited room for text-guided enhancement.
@@
We evaluated our model on RefCOCO/+/g [38], and the results are shown in Table 5. Combining the evaluation results from OmniLabel, we found that: L ED  +Qwen2-0.5B shows comparable performance to L ED  +InternLM2.5-7B despite 14×14$\times$14 × more parameters. This suggests a pattern of diminishing returns in scaling LLMs for visual grounding, supporting the hypothesis from [7] that small models can suffice for certain vision–language mappings.
Figure 5: OmniLabel prompts evaluated across LLM decoder layers 0024242424.
Table 6: OmniLabel performance (%) of L ED  based on InternVL2-1B (IVL2-1B) compare with SOTA GLIP-T–based data-generation pipelines.
Method AP AP
descr,categ descr
GLIP-T [15] 19.3 16.4
NEG-Text [45] 22.2 18.8
DesCo [14] 23.8 21.0
L ED  IVL2-1B 24.8 24.7
4.3 Comparison with synthetic-data baselines.
We integrate L ED  into the GLIP-T detector and evaluate on OmniLabel, juxtaposing it with two SOTA data-generation pipelines—NEG-Text [45] and DesCo [14]. Without synthesising any extra training data, L ED  attains the best overall score and the highest descr-AP.
As shown in the Table 6, L ED  achieves the highest overall score while requiring no synthetic labels, showcasing superior knowledge fusion over data-generation approaches. Notably, LED IVL2-1B particularly excels on the more challenging description samples, achieving a description AP of 24.7%, which is 3.7% higher than DesCo (21.0%), demonstrating its superior performance in complex semantic understanding and localization.
4.4 Deeper Decoder Layers Offer No Additional Benefit
To quantify the information contained at different decoder layers of the MLLM, we ablated adaptation prompts derived from each layer—including the pure ViT embeddings from the 00 “layer” (vision encoder only, without LLM)—and evaluated them on OmniLabel (see Figure 5). The ViT-only prompts yield minimal gains in the categorical AP, whereas prompts from early LLM decoder layers remain rich in semantic cues. In InternVL2-1B (24-layer Qwen2-0.5B), using only the first two decoder layers already yields significant improvements, with performance peaking at layer 4. Notably, deeper decoder layers do not guarantee better features. This aligns with our attention-map analysis in Section 3 and Figure 6(see Supplementary), which shows that image processing completes in early layers, “condensing” most visual information into those representations.
4.5 Detector with L ED  vs. Direct-LLM Substitution
As shown in the Appendix B(see Supplementary), replacing the detector’s visual embeddings with raw hidden states extracted from a frozen MLLM not only fails to match the original detector’s performance but can even cause a dramatic collapse in detection accuracy, confirming that latent features alone are neither geometrically nor semantically aligned with the detector’s visual space. Effective transfer therefore requires an explicit knowledge-fusion mechanism that jointly conditions on both language and vision, and the specialized detector remains indispensable for reliable performance in vision–language tasks.
@@ 5 Conclusions
Our work introduces a systematic framework that leveraging early MLLM decoder layers to boost visual grounding efficiency. We (1) demonstrate that shallow Transformer hidden states preserve rich spatial–semantic correlations; (2) design a zero-initialized cross-attention adapter for seamless MLLM-to-detector knowledge transfer; and (3) develop two adaptation-prompt schemes that sharpen descriptive precision while maintaining category accuracy. Integrating Qwen2-0.5B with Swin-Tiny yields a 3.82% OmniLabel gain at just 8.7% extra GFLOPs, while swapping in InternVL ViT pushes the improvement to 6.22%, alongside notable RefCOCO/+/g boosts. Extensive ablation studies across adapter variants, LLM scales, and decoder layers confirm our method’s robustness—especially for detailed and zero-shot queries. Our generalizable adapter design supports real-time vision–language applications and mitigates overfitting due to manual prompt bias, thereby advancing practical MLLM deployment.
Appendix A Experiment Details
A.1 Attention Maps in Qwen2-0.5B Decoding
Figure 6: Attention maps during the decoding process of Qwen2-0.5B.
Attention $\text{Attention Score}(Q,K)=\frac{QK^{T}}{\sqrt{d_{k}}}$ (10)
Where, $Q,K$ are the Query and Key matrix in the transformer Decoder layer, $\text{softmax}(\cdot)$as show in Figure 6, we have reached an approximate conclusion [2]. We can observe that in the first layer, attention is distributed relatively smoothly across different types of tokens. In the deeper layers, starting from local attention, attention scores are aggregated onto system prompts, instructions, and output tokens, while attention to image tokens becomes quite sparse. In the deeper layers, there are strong vertical lines (in the system prompts) that dominate most of the attention scores. The presence of these strong vertical lines indicates that certain input tokens remain highly attended to throughout the decoding process. If large-scale, high-intensity attention to image tokens were also maintained in the deeper layers, it would suggest that a significant amount of visual information is still needed for inference at those stages. However, as seen in the visualization, the deeper layers primarily focus on some key text tokens, which precisely indicates that the processing of images has already been completed in the earlier layers, and most of the image information has been "condensed" into the representations.
A.2 Adapter Architecture Detail
Figure 7: The adapter architecture $\mathrm{I}$ of our approach.
@@ A.2 Adapter Architecture Detail
A.3 Experiment Configuration
All experiments are conducted on eight NVIDIA A100 GPUs. During Stage 1 (pre-training) and Stage 2 (fine-tuning), we train the MLP layer for exactly one epoch each. For pre-training we use a batch size of 128, a learning rate of $1\times 10^{-3}$, weight decay of 0.03, and no warm-up. For fine-tuning, the warm-up ratio is set to 0.03, while the batch size, learning rate, and weight decay are fixed at 128, $4\times 10^{-5}$, and 0, respectively. In Stage 3 we train the adapter for one epoch with a batch size of 64, employing cosine learning-rate scheduling (base learning rate $2\times 10^{-4}$) without warm-up or weight decay. The MLP layer and the LLM are co-trained with the adapter using a separate learning rate of $4\times 10^{-5}$.
The hyper-parameters for Stages 1–3 are summarised in 7(a). To investigate the impact of model scale, we vary LLM parameter sizes as detailed in 7(b), with the results reported in Section 4.2.3. 7(c) lists the settings used to evaluate alternative adapter architectures; the corresponding OmniLabel scores appear in Table 3. Finally, 7(d) specifies the setup for examining different decoder layers, and the ablation results are presented in Section 4.4.
(a) Vision Decoder Share from Detector
@@ A.3 Experiment Configuration
$\ell_{LM}$ 2
$\mathbf{AP}$ to $\ell_{D}$ 6
Train Dataset Object365, COCO, Flickr30k
@@
Adapter arch. IV
(a) Experiment configuration details for vision decoder share from detector.
(b) Different MLLMs
@@
Train Dataset Object365, COCO, Flickr30k
adapter arch. IV
(b) Experiment configuration details for different LLMs.
(c) Adapter Architectures
@@
Train Dataset COCO, Flickr30k
(c) Experiment configuration details for different adapter architectures.
(d) Ablations: Decoder Layers
@@
Train Dataset Object365, COCO, Flickr30k
adapter arch. IV
(d) Ablation details for different decoder layers.
Table 7: Experiment configuration summary.
A.4 Dynamic Image Processing
Figure 9: Dynamic Image Processing
@@ A.4 Dynamic Image Processing
A.5 MLLM Alignment (Swin-T & Qwen2-0.5B)
@@
Appendix B Why MLLM Hidden States Cannot Replace Visual Detectors
One might be tempted to ask: “Can we simply replace a specialized visual detector with a LLM, MLLM or VLM and still achieve comparable performance?” To answer this question, Table 9 presents a head-to-head comparison between our detector-based pipeline and an end-to-end LLM approach across multiple benchmarks. As the results make clear, relying solely on the LLM without a dedicated object detector leads to a substantial drop in accuracy and consistency.
Table 9: Detector vs. MLLM Hidden-State Features on COCO-2017 val. Baselines are two GroundingDINO variants and vision-only SwinTiny-BERT; substitutions plug frozen decoder states from LLaVA-1.5 (L8) or InternVL-2B (L2/L8).
Method AP50:95 AP50 AP75
GroundingDINO-Pre 31.23 44.37 34.09
GroundingDINO-COCO SFT 57.23 73.27 63.18
SwinTiny-BERT 41.97 57.45 45.85
LLaVA-L8 24.19 39.47 24.94
IVL2-L2 38.28 55.25 41.18
IVL2-L8 42.64 63.77 44.92
GroundingDINO-Pre establishes a detector base on Open-GroundingDINO [19] that has never been exposed to COCO, being trained only on Objects365, GoldG, and Cap4M; GroundingDINO-COCO SFT fine-tunes the same weights on COCO SwinTiny-BERT serves as a vision-only control whose Swin-Tiny backbone and BERT text branch use only same framework as Open-GroundingDINO [19] is identical to that used in the multi-modal runs yet, like the MLLMs, has never seen COCO or any object-detection data. The three substitution variants—LLaVA-L8, IVL2-L2, and IVL2-L8—replace the Swin feature map with hidden states taken from decoder layers 2 or 8 of LLaVA-1.5 or InternVL-2B, enabling a direct test of whether raw MLLM representations can stand in for detector-oriented visual embeddings.
As show int the Table 9, fine-tuning the vision backbone on COCO (GDINO-COCO) improves AP50:95 by +26 points over its out-of-domain counterpart, underscoring the importance of spatial alignment learned from explicit detection supervision. By contrast, treating language–vision model (MLLM) activations as a drop-in surrogate for visual tokens is ineffective: the strongest variant (IVL2-L8) still trails the fully visual SwinTiny-BERT by $-0.67$ AP50:95 and remains $-14.6$ points below the COCO-tuned baseline. The gap widens to $-33$ points when substituting LLaVA-L8, revealing that decoder-layer semantics alone carry minimal localisation cues.
This discrepancy arises because MLLM hidden states are produced after heavy token mixing and rotary positional encoding optimised for caption generation; they lack the multi-scale geometry and inductive biases embedded in convolutional or ViT backbones. Without an explicit fusion mechanism that jointly conditions on language and vision, the detector cannot recover accurate object coordinates, leading to systematic localisation failure. The comparison between SwinTiny-BERT (whose Swin weights, like the MLLMs, have never seen COCO) and the substitution runs highlights the central claim of this study: knowledge fusion is indispensable—naive replacement of visual embeddings with isolated MLLM features is not a viable route to open-vocabulary detection.
Appendix C Case Study
Notably, the adaptation prompt functions as expected, guiding the detector rather than allowing the adapter to dominate the grounding task. As seen in the recognition results in Figure 10, GroundingDINO often fails to locate individuals, instead detecting entire objects, as in the case of "The computer monitors on the desk that are turned on." Additionally, it struggles to interpret positional relationships in text descriptions, such as "the sandwiches that are each on the same plate with the other." With the adaptation prompt, the detector retains an overall understanding while incorporating individual positional relationships. Furthermore, it corrects errors in spatial logic relationships.
Appendix D Implementation Details
D.1 Dataset Introduction
Table 10 provides the attributes, sizes, and functional descriptions of all datasets used in our stages 1, 2, 3 and evaluation, to illustrate the support they offer during the process. All datasets are distributed under licences that permit non-commercial academic research.